# ON THE AVERAGE ULTRAVIOLET EMISSION LINE SPECTRA OF HIGH-REDSHIFT GALAXIES: HOT AND COLD, CARBON-POOR, NITROGEN-MODEST, AND OOZING IONIZING PHOTONS

### v1

@@ 4.1.2 Forbidden Oxygen Lines
We compute the electron temperature in the high ionization zone from the [O iii]1666/5008 Å ratio using version 1.1.15 of PyNeb (Luridiana et al., 2015), with tranistion probabilities from Froese Fischer & Tachiev (2004) and collision strengths from Aggarwal & Keenan (1999) and Kisielius et al. (2009). We again correct all emission lines for dust obscuration using Reddy et al. (2015, 2016) attenuation law, using dust reddenings measured from the H/H line ratio assuming an intrinsic value of 2.86, which is appropriate for K gas. We note here that the final O++ temperatures are almost twice this temperature, which would decrease the intrinsic H/H ratio to 2.76, and require more attenuation to match the observed ratios. However, as we show in Section 4.1.1, the average temperatures in the nebula are more consistent with K when derived from radiation produced by recombining hydrogen, so we do not modify this intrinsic line Balmer decrement.
Because of the low resolution of NIRSpec/PRISM, we do not resolve any of the density-dependent close doublets, so assume an electron density () of 250 cm-3 throughout for consistency with other studies at similar redshifts (Sanders et al., 2018; Reddy et al., 2023; Revalski et al., 2024). This provides the temperatures in the O++ zone (hereafter ), which represents only a fraction of the nebula, and does not sample the likely-dominant regions where oxygen is only once-ionized. We do not have access to the very weak auroral lines of [O ii] at 7320,7330Å, and do not detect the [N ii]Å line to measure in the less ionized medium, so instead use the scaling function from Garnett et al. (1995) to estimate the temperature of the O+ zone (hereafter ).
@@ 4.2 Metal Abundances
The derivation of these temperatures points directly towards the abundances of O, C, and N using the temperature-dependent collisional and recombination rate coefficients calculated in PyNeb (Luridiana et al., 2015). We now go through each of these elements in turn.
@@
Oxygen abundances are derived using the dust corrected ratio of [O ii]3727/H and to obtain the abundance O+/H+. With the O+ temperatures derived above we fix the recombination and collisional coefficients for H and O, and calculate directly the abundance that would produce the observed line ratio. Analogously we take [O iii]5008/H and to obtain the abundance O++/H+. We sum O+/H+ + O++/H+ to obtain the total oxygen abundance O/H. For our standard set of assumptions, the nebular oxygen abundance falls in the range to 8.0 (Table 7).
@@
Carbon abundances are derived using the dust corrected ratio of C iii]1909/H and to obtain the abundance C++/H+. Ionization correction factors are calculated following Berg et al. (2019b) by first deriving the ionization parameter, , from the O32 ratio. For galaxies showing C iv in emission we also tested the summation of the C iv and C iii] abundances to avoid the need for ICFs. We report these measurements in Table 7 and show the variation of the (C/O) ratio with in Figure 8. For our standard set of assumptions, the carbon/oxygen ratio covers very little range, and all the five stacks show ratios consistent with , regardless of O/H.
@@
Nitrogen abundances are derived using the dust corrected ratio of N iii]1749/H and N iv]1483,86/H, together with to obtain the abundance N++/H+ and N3+/H+, respectively. We use fluxes derived from numerical integration of the emission lines. To bypass the need for ICFs, we calculate N/O using (N+++N3+)/O++ or N++/O++ when N iv] was not detected. We report these measurements in Table 7 and show the variation of the (N/O) ratio with in Figure 9, using our standard assumptions in the left-hand panel. While uncertainties are large, the measurements in our stacks lie between the high- galaxies and estimates from local dwarf galaxies and extragalactic H ii regions.
@@ 4.3 Tests of Fundamental Assumptions
Without access to density-sensitive doublet like [O ii], we had to assume , taking a value of 250 cm-3 for consistency with other high- measurements (e.g. Sanders et al., 2023, and others). The UV lines all have critical densities in excess of cm-3, but the optical lines of [O ii]3727 and [O iii]5008 used for both oxygen abundances and temperature measurements do not. We therefore run all of our calculations with a high value of cm-3.
This intervention has two main results. The first is on the electron temperatures, for which which decreases systematically by %. In fact, when calculating the using the Garnett et al. (1995) formulation, the [O iii] temperatures are only % hotter than those derived by the Balmer jump. As a consequence of this, all abundances increase: increases by the substantial increment of 0.5 dex, shifting all the points in Figure 8 to the right. Abundances of C and N, however, change by similar factors and the (C/O) and (N/O) ratios are close to constant (right panel of Figure 9. This shifts our points in the (N/O) diagram to become more in-line with the local H ii regions, suggesting no enhancement of nitrogen. However, the individual high- galaxy estimates would shift by comparable amounts if the density were so high, implying much smaller levels of N enhancement than otherwise expected.
@@ 5 Results III: Production and Escape of Ionizing Radiation
We make several inferences on the production and escape of ionizing radiation, and compute two of the main quantities in the ionizing emissivity: the ionizing photon production efficiency (), and the escape fraction of ionizing radiation (). is defined as the number of ionizing photons produced per unit time per unit UV luminosity: [Hz erg-1], where is the number of ionizing photons produced per second for is the luminosity density per unit frequency (e.g. Robertson et al., 2015, but see Tarrab (1985) and Mas-Hesse & Kunth (1991) for some historical perspectives on as the ‘’ and ‘-parameter’). We compute from the optical emission line spectrum using , where is the dust corrected H luminosity in erg s-1. is computed from the H and H fluxes assuming the Reddy et al. (2015) attenuation law. We show the resulting calculation of , projected against the variable that defines the stacks, in the upper left panel of Figure 10. log(/[Hz erg-1]) ranges from 25.15 to 25.65, increasing monotonically with , which is expected as both EWs and encode line emission divided by stellar continuum. Values of are reported in Table 8.
@@
Naturally the galaxies in our sample lie at redshifts that are far too high to measure the output of ionizing continuum, so for we must rely upon independent diagnostics. For this we turn to the escape fraction of Mg ii radiation ( Henry et al., 2018; Chisholm et al., 2020; Xu et al., 2022, 2023) and the so-called [S ii]-deficit (Alexandroff et al., 2015; Wang et al., 2019, 2021). Xu et al. (2022, see also ) have published polynomial fitting functions to describe the relationship between the ratio of both components of the Mg ii doublet ( and 2803 Å, independently) to the [O iii]5008 luminosity, compared to the O32 ratio for both density and ionization bounded nebulae, based upon extensive sets of CLOUDY simulations. We adopt these fits to compute the expected luminosities in the Mg ii doublet using our [O ii] and [O iii] measurements, and sum them to obtain the expected intrinsic Mg ii. We then obtain by dividing the measured Mg ii luminosity by the calculated value. We show the result in the upper right panel of Figure 10. We make no attempt to correct for interstellar absorption of the resonant Mg ii line which, unlike in the high-resolution ground-based observations of Xu et al. (2023) we obviously cannot resolve. We instead plot negative as open symbols. ranges from below 0 to almost 100%. Values of are reported in Table 8.
@@
For the [S ii] diagnostics we follow a similar procedure, and calculate both the [O iii]5008/H and [S ii]6717,31/H ratios directly in the stacked spectra. We then compute the [S ii]-deficit ([S ii]) using the formulae in Wang et al. (2021), where [S ii] is defined as the distance of a galaxy’s measured log([S ii]/H) to the expected value based upon its log([O iii]/H) and the locus of SDSS points on the BPT diagram. We plot these ratios in the lower left panel of Figure 10, and report values of [S ii] in Table 8. log([O iii]/H) are relatively constant at , but log([S ii]/H) drifts monotonically away from the SDSS locus for increasing . Estimates of [S ii] are given in Table 8.
@@
We finally compute the O32 ratio directly from the line luminosities, and UV continuum slope, , from the stacks with luminosity normalized at 3650 Å. We show these in the lower right panel of Figure 10. Across the sequence of increasing , UV slopes become bluer and O32 values higher. We illustrate the LyC-leaking region of the range from Chisholm et al. (2022) without shading. has already been reported in Table 1.
@@
Throughout Figure 10 we contrast results from these stacks with individual measurements made in the Low-Redshift Lyman Continuum Survey, using data from Saldana-Lopez et al. (2022) for , Chisholm et al. (2022) for , Flury et al. (2022b) for , Xu et al. (2023) for Mg ii, and Wang et al. (2021) for [S ii].
@@ 6.1 UV Line Strengths and Implications for the Ionizing Sources
Our stacks reveal an increase in the equivalent widths of rest UV lines (e.g. C iii]) with the EW of [O iii] in the optical (Figures 2 and 3 and Section 3.1). This is a very natural result, especially given the increase in , shown in Figure 10 which normalizes line flux by UV continuum directly. Similar phenomena have been seen before (Berg et al., 2016; Senchyna et al., 2017) and our results for C iii] agree well with those of Ravindranath et al. (2020), who obtained and compiled a large sample of galaxies over a range of redshifts. This correlation was also the probable explanation for the high detection rate of C iii] (Stark et al., 2017; Laporte et al., 2017) in Spitzer-selected [O iii]-emitters at (Roberts-Borsani et al., 2016). Large numbers of C iii]-bright galaxies have also been identified at mid-, most notably among the sample of 450 Main-Sequence galaxies studied by Nakajima et al. (2018); Llerena et al. (2022) at . These show an average of just 2.2 Å, which is consistent only with the least intensely star-forming subsets of our sample. This value of is already doubled by the 2nd bin in the stacks, and exceeded by a factor of in our most intensely star-forming stack, which shows Å. Nakajima et al’s models suggest an upper limit to of 12 Å for a zero-age starburst comprised of ‘conventional’ stars, which is exceeded here by %. We stress again that this stacked spectrum of the highest galaxies is the average of 130 objects, and it is surprising that the average source should lie % above this limit in if conditions in the two galaxy samples were comparable.
@@
Notably, the factor of increase in in our stacks is significantly larger than the factor of increase in that happens across the same range (Figure 10). This implies that something more than a simple increase in must be at play. For example, even spectra with more evolved stellar population (showing a normal stellar Balmer break, with ages Myr) also show estimates as high as 25.2. We also see an increase in the ratio of the C iv/C iii] fluxes (Figure 4), which may point towards and evolution of the shape of the ionizing continuum and/or other nebular conditions (e.g. a decrease in temperature). Indeed, Nakajima et al. (2018) have showed that the strength of C iii] and C iv peaks at different metallicities in their photoionization models, as a result of changes in the gas temperature and the ionization structure of the ISM. A number of possibilities may underlie the trends seen in the stacks: (1.) lower gas metallicities may be implicated in the upward evolution of the C iii] EW (e.g. Ravindranath et al., 2020; Roberts-Borsani et al., 2024) – indeed we do find a general decrease in oxygen abundance across the stacks, ranging from 12+log(O/H) to 8.0, which covers exactly the range where C iii] EW should peak. These UV lines are also extremely temperature sensitive and are thus mainly visible in low-metallicity targets: metallicity seemed the factor driving most their presence/absence in CLASSY (Mingozzi et al., 2024). (2.) AGN may play a more dominant but currently unseen role, and their harder ionizing spectra produce more eV photons than typical stars. However we do not see other trends typical of AGN, and the fact that UV luminosities decrease and is consistent with ordinary star forming galaxies seems to disfavor the AGN interpretation. (3.) the higher galaxies may be the result of galaxies hosting a higher fraction of common-envelope binary stars and/or stripped staras are present in these higher stacks. This would provide qualitatively similar changes to the ionizing spectra (Götberg et al., 2019) as for AGN over the comparatively narrow energy range of 13.6 to 24 eV. (4.) low-velocity shocks become more prevalent (e.g. Jaskot & Ravindranath, 2016), increasing the strength of semi-forbidden C iii] compared to the continuum.
@@
Interestingly, the behavior of the C iv emission is not nearly as extreme: it peaks at only Å. This is the strength of the C iv-emitter sample of VANDELS galaxies presented in (Mascia et al., 2023), which have a median of 7.6 Å. We note, of course, that statics on a C iv-emission only sample will be biased high by excluding the dominant fraction of galaxies that do not have C iv fluxes above the detection threshold. However, while the C iii] EWs of our stacks are extreme by the standards of mid- individual galaxies, the C iv emission is not.
@@ 6.2 Gas Temperatures in Early Galaxies
Electron temperatures derived from the ratio of [O iii]5008/1666Å emission lines fall in the range 14,000 to 21,000 Kelvin and incrase monotonically with the intensity of star formation within the stack: this change occurs as ages decrease from Myr to just 3 Myr (right panel of Figure 6). This trend in temperature is expected as episodes of star formation become younger and contain a higher fraction of more massive stars (although see the caveats concerning AGN contamination above): ionization by stars that are hotter on average transfer more kinetic energy to the free electrons in the photoionization processes. At the same time, cooling is less efficient at lower metallicities, and we also see a clear change in the nebular oxygen abundance, also decreasing as temperature increase (and ages decrease), as shown in Figure 8.
@@
Are the oxygen temperatures too high? One possibility is that, when using the [O iii]1666/5008Å ratio, small uncertainties in the dust attenuation lead to large over-corrections and O++ temperatures that are too high. However, we believe this not to be the case for at least three reasons. Firstly, we can also measure in three of the stacks using the [O iii]4363/5008Å ratio: this is not nearly as sensitive to attenuation law variations, and yields temperatures consistent with those from [O iii]1666/5008Å. We also have He ii measurements at both 1640Å (Balmer alpha) and 4687Å (Paschen alpha), which are very close in wavelength to the 1666 and 5008Å [O iii] lines. If we instead take the ratio of [O iii]/He ii ratios (O iii]1666/He ii1640)/([O iii]5008/He ii4687) any effect of dust attenuation is canceled. Thus we believe over-correction for dust is not responsible for the high temperatures.
@@
We have also assumed a fixed electron density of cm-3 throughout these calculations, and in principle the [O iii]1666/5008Å ratios could be artificially high if the 5008Å line is collisionally de-excited. However, the critical density of the O++ 1D2 level is cm-3, which is far higher than expected for the ISM density of the average galaxy.
@@
Our estimates by both methods also broadly agree with other measurements made at similar redshifts and in analog galaxies. Schaerer et al. (2022a) find K from [O iii]4363 for the two galaxies, Trump et al. (2023) obtain ,000 K from [O iii] at , and Curti et al. (2023) find some temperatures in excess of 25,000 K. Nakajima et al. (2023) carefully re-analyze several previously published samples of high-, NIRSpec-derived measurements, finding many have been over-estimated, but still report temperatures ,000 K. Similar numbers and ranges are reported by Sanders et al. (2024), which are all consistent with our [O iii] results in samples of what should be the most highly ionized, hottest starbursts. In the low- analog starburst galaxies, Mingozzi et al. (2022) find temperatures in the range 12,000 to 20,000 K using the UV lines for the CLASSY sample, while temperatures in excess of 19,000 K are also reported for LzLCS using the optical lines (Flury et al., 2022b). Nothing about our [O iii] temperature is surprising.
@@
Few examples of Balmer jump temperatures exist at comparable redshifts. Welch et al. (2024) measured Balmer Jump temperatures of 12,000 K in the dominant LyC-emitting cluster of the Sunburst arc (Rivera-Thorsen et al., 2019), which is consistent with our BJ estimates for the strongest [O iii]-emitters. We note that they also measured [O iii] optical temperatures that were not as extreme as we found (i.e. the discrepancy is smaller) but we also point out that this cluster is thought to be extreme, showing strong nitrogen enrichment, very high densities, and significant UV emission lines (Meštrić et al., 2023; Pascale et al., 2023, see below). On the other hand our Balmer Jump temperatures are in stark disagreement with those recently reported by Katz et al. (2024), who find temperatures of 20,000 K and sometimes much higher when using a similar approach, and whose sample must contain some of the same objects as ours. It is plausible that our stacking of hundreds of galaxies picks out a more representative and less extreme sub-sample, with lower average than those of Katz et al. (2024). While not derived from the Balmer Jump, very interesting estimate of the O+ zone temperature has was reported by Sanders et al. (2023). They measured of 12,000 and 9000 K using using auroral lines of [O ii] in 2 galaxies observed with JWST at , where the line formation should occur in the same zone as our recombination continuum based estimates.
@@
Temperature fluctuations are invoked to explain abundance discrepancies in planetary nebulae and H ii regions (comparing collisional metal lines with recombination metal lines, in the very local Universe only; e.g. Méndez-Delgado et al. 2023). The gas is clumpy and has uneven illumination, and the distribution of temperatures is set as the cooling rate varies. Only the hotter parts of the nebula contribute to the production of auroral line radiation, skewing the temperature upwards (and the oxygen abundance downwards). One may expect similar behavior in the integrated spectra of galaxies, but this was not the result found by Guseva et al. (2006). These authors looked at local blue compact galaxies, finding good agreement between [O iii]4363/5008 and Balmer (and Paschen) Jump temperature estimates. It is plausible, of course, that some of this result could be explained by aperture effects: if the brightest optical point sources are observed with narrow slits the spectra are likely biased towards the hottest regions of the nebulae, revealing high temperature that are also matched by the Balmer Jump. However, the larger, cooler, lower ionized zones that still contribute to the free bound continuum would not be captured by small slits and large discrepancies could be avoided. Our spectroscopic apertures, which have sizes of several kpc, would capture the full extent of the nebulae where temperatures are genuinely lower than in the small sub-regions that dominate the [O iii] emission.
6.3 Metal Abundances

### v2

@@ 4.1.2 Forbidden Oxygen Lines
We compute the electron temperature in the high ionization zone from the (dust-corrected) [O iii]5008/1666 Å ratio, using version 1.1.15 of PyNeb (Luridiana et al., 2015), with transition probabilities from Froese Fischer & Tachiev (2004) and collision strengths from Aggarwal & Keenan (1999) and Kisielius et al. (2009). We note here that the final O++ temperatures are almost twice the assume value for the intrinsic Balmer decrement, which would decrease the intrinsic H/H ratio from 2.86 to 2.76, and require more attenuation to match the observed ratios. However, as we show in Section 4.1.1, the average temperatures in the nebula are more consistent with K when derived from radiation produced by recombining hydrogen, so we do not modify this intrinsic line Balmer decrement.
Because of the low resolution of NIRSpec/PRISM, we do not resolve any of the density-dependent close doublets, so assume an electron density () of 250 cm-3 throughout for consistency with other studies at similar redshifts (Sanders et al., 2018; Reddy et al., 2023; Revalski et al., 2024). This provides the temperatures in the O++ zone (hereafter ), which represent only a fraction of the nebula, and do not sample the likely-dominant regions where oxygen is only once-ionized. We do not have access to the very weak auroral lines of [O ii] at 7320,7330Å, and do not detect the [N ii]Å line to measure in the less ionized medium, so instead use the scaling function from Garnett et al. (1995) to estimate the temperature of the O+ zone (hereafter ).
@@ 4.2 Metal Abundances
The derivation of these temperatures leads us directly towards the abundances of O, C, and N, using the temperature-dependent collisional and recombination rate coefficients calculated in PyNeb (Luridiana et al., 2015). We now go through each of these elements in turn.
@@
Oxygen abundances are derived using the ratio of [O ii]3727/H and to obtain the abundance O+/H+. With the O+ temperatures derived above we fix the recombination and collisional coefficients for H and O, and calculate directly the abundance that would produce the observed line ratio. By analogy, we take [O iii]5008/H and to obtain the abundance O++/H+. We sum O+/H+ + O++/H+ to obtain the total oxygen abundance O/H. For our standard set of assumptions, the nebular oxygen abundance falls in the range to 8.0 (see Table 8).
@@
Carbon abundances are derived using the ratio of C iii]1909/H and to obtain the abundance C++/H+. Ionization correction factors are calculated following Berg et al. (2019b) by first deriving the ionization parameter, , from the O32 ratio. For galaxies showing C iv in emission we also tested the summation of the C iv and C iii] abundances to avoid the need for ICFs. We report these measurements in Table 8 and show the variation of the (C/O) ratio with in Figure 8. For our standard set of assumptions, the carbon/oxygen ratio covers very little range, and all the five stacks show ratios consistent with , regardless of O/H.
@@
Nitrogen abundances are derived using the ratio of N iii]1749/H and N iv]1483,86/H, together with to obtain the abundance N++/H+ and N3+/H+, respectively. We use fluxes derived from numerical integration of the emission lines. To bypass the need for ICFs, we calculate N/O using (N+++N3+)/O++ or N++/O++ when N iv] was not detected. We report these measurements in Table 8 and show the variation of the (N/O) ratio with in Figure 9, using our standard assumptions in the left-hand panel. While uncertainties are large, the measurements in our stacks lie between the high- galaxies and estimates from local dwarf galaxies and extragalactic H ii regions.
@@ 4.3 Tests of Fundamental Assumptions
Without access to density-sensitive doublet like [O ii], we had to assume , taking a value of 250 cm-3 for consistency with other high- measurements (e.g. Sanders et al., 2023). The UV lines all have critical densities in excess of cm-3, but the optical lines of [O ii]3727 and [O iii]5008 used for both oxygen abundances and temperature measurements do not. We therefore run all of our calculations with a high value of cm-3.
This intervention has two main results. The first is on the electron temperatures, for which decreases systematically by %. In fact, when calculating using the Garnett et al. (1995) formulation, the [O iii] temperatures become only % hotter than those derived by the Balmer jump. As a consequence of this, all abundances increase: increases by the substantial increment of 0.5 dex, shifting all the points in Figures 8 and 9 to the right. Abundances of C and N, however, change by similar factors and the (C/O) and (N/O) ratios are close to constant (right panel of Figure 9). This shifts our points in the (N/O) diagram to become more in-line with the local H ii regions, suggesting no enhancement of nitrogen. Note, of course, that the individual high- galaxy estimates would shift by comparable amounts if the density were so high, implying much smaller levels of N enhancement than otherwise have been reported.
@@ 5 Results III: Production and Escape of Ionizing Radiation
We make several inferences on the production and escape of ionizing radiation, and compute two of the main quantities in the ionizing emissivity: the ionizing photon production efficiency (), and the escape fraction of ionizing radiation (). is defined as the number of ionizing photons produced per unit time per unit UV luminosity: [Hz erg-1], where is the number of ionizing photons produced per second for is the luminosity density per unit frequency (e.g. Robertson et al., 2015, but see Tarrab (1985) and Mas-Hesse & Kunth (1991) for some interesting historical perspectives on as the ‘’ and ‘-parameter’). We compute from the optical emission line spectrum using , where is the dust corrected H luminosity in erg s-1. is computed from the H and H fluxes assuming the Reddy et al. (2015) attenuation law. We show the resulting calculation of , projected against the variable that defines the stacks, in the upper left panel of Figure 10. log(/[Hz erg-1]) ranges from 25.15 to 25.65, and increases monotonically with – this relation has been seen numerous times in the past (Chevallard et al., 2018; Tang et al., 2019; Simmonds et al., 2024a), and is expected since both EWs and encode line fluxes divided by stellar continuum. Values of are reported in Table 9.
@@
Naturally the galaxies in our sample lie at redshifts that are far too high to measure the output of ionizing continuum, so for we must rely upon independent diagnostics. For this we turn to the escape fraction of Mg ii radiation ( Henry et al., 2018; Chisholm et al., 2020; Xu et al., 2022, 2023) and the so-called [S ii]-deficit (Alexandroff et al., 2015; Wang et al., 2019, 2021). Xu et al. (2022, see also ) have published polynomial fitting functions to describe the relationship between the ratio of both components of the Mg ii doublet ( and 2803 Å, independently) to the [O iii]5008 luminosity, compared to the O32 ratio for both density and ionization bounded nebulae, based upon extensive sets of CLOUDY simulations. We adopt these fits to compute the expected luminosities in the Mg ii doublet using our [O ii] and [O iii] measurements, and sum them to obtain the expected intrinsic Mg ii. We then obtain by dividing the measured Mg ii luminosity by the estimated intrinsic value. We show the result in the upper right panel of Figure 10. We make no attempt to correct for interstellar absorption of the resonant Mg ii line which, unlike in the high-resolution ground-based observations of Xu et al. (2023) we obviously cannot resolve (the effect would be to increase our estimates). We instead plot negative as open symbols. ranges from below 0 to almost 100%. Values of are reported in Table 9.
@@
For the [S ii] diagnostics we follow a similar procedure, and calculate both the [O iii]5008/H and [S ii]6717,31/H ratios directly in the stacked spectra. We then compute the [S ii]-deficit ([S ii]) using the formulae in Wang et al. (2021), where [S ii] is defined as the distance of a galaxy’s measured log([S ii]/H) from the expected value based upon its log([O iii]/H) and the locus of SDSS points on the BPT diagram. We plot these ratios in the lower left panel of Figure 10, and report values of [S ii] in Table 9. log([O iii]/H) are relatively constant at , but log([S ii]/H) drifts monotonically away from the SDSS locus for increasing . Estimates of [S ii] are given in Table 9.
@@
We finally compute the O32 ratio directly from the line luminosities, and UV continuum slope, , from the stacks with luminosity normalized at 3650 Å (second left panel of Figure 2). We show these in the lower right panel of Figure 10. Across the sequence of increasing , UV slopes become bluer and O32 values higher. We illustrate the LyC-leaking region of the range from Chisholm et al. (2022) without shading. has already been reported in Table 1.
@@
Throughout Figure 10 we contrast results from these stacks with measurements made in individual galaxies from the Low-Redshift Lyman Continuum Survey, using data from Saldana-Lopez et al. (2022) for , Chisholm et al. (2022) for , Flury et al. (2022b) for , Xu et al. (2023) for Mg ii, and Wang et al. (2021) for [S ii].
@@ 6.1 UV Line Strengths and Implications for the Ionizing Sources
Our stacks reveal an increase in the equivalent widths of rest UV lines (e.g. C iii]) with the EW of [O iii] in the optical (Figures 2 and 3 and Section 3.1). This is a very natural result, especially given the increase in , shown in Figure 10 which normalizes line flux by UV continuum directly. Similar phenomena have been seen before (Berg et al., 2016; Senchyna et al., 2017) and our results for C iii] agree well with those of Ravindranath et al. (2020), who obtained and compiled a large sample of galaxies over a range of redshifts. This correlation was also the probable explanation for the high detection rate of C iii] (Stark et al., 2017; Laporte et al., 2017) in Spitzer-selected [O iii]-emitters at (Roberts-Borsani et al., 2016).
Large numbers of C iii]-bright galaxies have also been identified at mid-, most notably among the sample of 450 Main-Sequence galaxies studied by Nakajima et al. (2018) and Llerena et al. (2022) at . These show an average of just 2.2 Å, which is consistent only with the least intensely star-forming subsets of our sample. This value of is already doubled by the 2nd bin in our stacks, and exceeded by a factor of in our most intensely star-forming stack, which shows Å. Nakajima et al’s models suggest an upper limit to of 12 Å for a zero-age starburst comprised of ‘conventional’ stars, which is exceeded here by %. We stress again that this stacked spectrum of the highest galaxies is the average of 130 objects, and it is surprising that the average source should lie % above this limit in if conditions in the two galaxy samples were comparable.
@@
Notably, the factor of increase in over our range is significantly larger than the factor of increase in (Figure 10). This implies that something more than a simple increase in must be at play. For example, even spectra with more evolved stellar population (showing a normal stellar Balmer break, with ages Myr) also show estimates as high as 25.2. We also see an increase in the ratio of the C iv/C iii] fluxes (Figure 4), which may point towards and evolution of the shape of the ionizing continuum and/or other nebular conditions (e.g. an increase in temperature associated with lower gas-phase metallicity). Indeed, Nakajima et al. (2018) have showed that the strength of C iii] and C iv peaks at different metallicities in their photoionization models, as a result of changes in the gas temperature and the ionization structure of the ISM. A number of possibilities may underlie the trends seen in the stacks: (1.) lower gas metallicities may be implicated in the upward evolution of the C iii] EW (e.g. Ravindranath et al., 2020; Roberts-Borsani et al., 2024) – indeed we do find a general decrease in oxygen abundance across the stacks, ranging from 12+log(O/H) to 7.5, which covers exactly the range where C iii] EW should peak. These UV lines are also extremely temperature sensitive and are thus mainly visible in low-metallicity targets: metallicity seemed the factor driving most their presence/absence in CLASSY (Mingozzi et al., 2024). (2.) AGN may play a more dominant but currently unseen role, and their harder ionizing spectra produce more eV photons than typical stars. It is interesting in this context that the [C ii]2326 feature becomes visible only in the highest stack. However we do not see other trends typical of AGN, and the fact that UV luminosities decrease and is consistent with ordinary star forming galaxies may disfavor the AGN interpretation. (3.) the higher galaxies may be the result of galaxies hosting a higher fraction of common-envelope binary stars and/or stripped stars are present in these higher stacks. This would provide qualitatively similar changes to the ionizing spectra (Götberg et al., 2019) as for AGN over the comparatively narrow energy range of 13.6 to 24 eV. (4.) low-velocity shocks become more prevalent (e.g. Jaskot & Ravindranath, 2016), increasing the strength of semi-forbidden C iii] compared to the continuum.
@@
Interestingly, the behavior of the C iv emission is not nearly as extreme: it peaks at only Å. This is the strength of the C iv-emitter sample of VANDELS galaxies presented in (Mascia et al., 2023), which have a median of 7.6 Å. We note, of course, that statistics taken on a C iv-emission only sample will be biased high by excluding the dominant fraction of galaxies that do not have C iv fluxes above the detection threshold. We finally remake that, while the C iii] EWs of our stacks are extreme by the standards of mid- individual galaxies, the C iv emission appears not to be.
@@ 6.2 Gas Temperatures in Early Galaxies
Electron temperatures derived from the ratio of [O iii]5008/1666Å emission lines fall in the range 14,000 to 21,000 Kelvin and increase monotonically with the intensity of star formation within galaxies: this change occurs as ages decrease from Myr to just 3 Myr (right panel of Figure 6). This trend in temperature is expected as episodes of star formation become younger and contain a higher fraction of more massive stars (although see the caveats concerning AGN contamination above): ionization by stars that are hotter on average transfer more kinetic energy to the free electrons in the photoionization processes. At the same time, cooling is less efficient at lower metallicities, and we also see a clear change in the nebular oxygen abundance, also decreasing as temperature increase (and ages decrease), as shown in Figure 8.
@@
Are the oxygen temperatures too high? One possibility is that, when using the [O iii]5008/1666 Å ratio, small uncertainties in the dust attenuation lead to large over-corrections and O++ temperatures that are too high. However, we believe this not to be the case for at least three reasons. Firstly, we can also measure in three of the stacks using the [O iii]5008/4363 Å ratio: this is not nearly as sensitive to attenuation law variations, and yields temperatures consistent with those from [O iii]5008/1666Å. We also have He ii measurements at both 1640Å (Balmer alpha) and 4687Å (Paschen alpha), which are very close in wavelength to the 1666 and 5008Å [O iii] lines. If we instead take the ratio of [O iii]/He ii ratios (O iii]1666/He ii1640)/([O iii]5008/He ii4687) any effect of dust attenuation is canceled. Thus we believe over-correction for dust is not responsible for the high temperatures.
@@
We have also assumed a fixed electron density of cm-3 throughout these calculations, and in principle the [O iii]5008/1666 Å ratios could be artificially high if the 5008Å line is collisionally de-excited. However, the critical density of the O++ 1D2 level is cm-3, which is far higher than expected for the ISM density of the average galaxy.
@@
Our estimates by both methods also broadly agree with other measurements made at similar redshifts and in analog galaxies. Schaerer et al. (2022a) find K from [O iii]4363 for the two galaxies, Trump et al. (2023) obtain ,000 K from [O iii] at , and Curti et al. (2023) find some temperatures in excess of 25,000 K. Nakajima et al. (2023) carefully re-analyze several previously published samples of high-, NIRSpec-derived measurements, finding many have been over-estimated, but still report temperatures as high as ,000 K for a number of systems. Similar numbers and ranges are reported by Sanders et al. (2024), which are all consistent with our [O iii] results in samples of what should be the most highly ionized, hottest starbursts. In the low- analog starburst galaxies, Mingozzi et al. (2022) find temperatures in the range 12,000 to 20,000 K using the UV lines for the CLASSY sample, while temperatures in excess of 19,000 K are also reported for LzLCS using the optical lines (Flury et al., 2022b). Nothing about our [O iii] temperature is surprising.
@@
Few examples of Balmer jump temperatures exist at comparable redshifts. Welch et al. (2024) measured Balmer Jump temperatures of 12,000 K in the dominant LyC-emitting cluster of the Sunburst arc (Rivera-Thorsen et al., 2019), which is consistent with our BJ estimates for the strongest [O iii]-emitters. We note that they also measured [O iii] optical temperatures that were not as extreme as we found (i.e. the discrepancy is smaller) but we also point out that this cluster is thought to be extreme, showing strong nitrogen enrichment, very high densities, and significant UV emission lines (Meštrić et al., 2023; Pascale et al., 2023, see below). On the other hand our Balmer Jump temperatures are in stark disagreement with those recently reported by Katz et al. (2024), who find temperatures of 20,000 K and sometimes much higher when using a similar approach, and whose sample must contain some of the same objects as ours. It is plausible that our stacking of hundreds of galaxies picks out a more representative and less extreme sub-sample, with lower average than those of Katz et al. (2024). While not derived from the Balmer Jump, very interesting estimate of the O+ zone temperature has was reported by Sanders et al. (2023). They measured of 12,000 and 9000 K using using auroral lines of [O ii] in 2 galaxies at , where the line formation should occur in the same zone as our recombination continuum based estimates.
@@
Temperature fluctuations are invoked to explain abundance discrepancies in planetary nebulae and H ii regions (comparing collisional metal lines with recombination metal lines, in the very local Universe only; E.g. Méndez-Delgado et al. 2023). If the gas is clumpy and has uneven illumination, then the distribution of temperatures will be set as the cooling rate varies. Only the hotter parts of the nebula contribute to the production of auroral line radiation, skewing the temperature upwards (and the inferred oxygen abundance downwards). One may expect similar behavior in the integrated spectra of galaxies, but this was not the result found by Guseva et al. (2006). These authors looked at local blue compact galaxies, finding good agreement between [O iii]5008/4363 and Balmer (and Paschen) Jump temperature estimates. It is plausible, of course, that some of the Guseva et al. (2006) result could be explained by aperture effects: if the brightest optical point sources are observed with narrow slits the spectra are likely biased towards the hottest regions of the nebulae, revealing high temperature that are also matched by the Balmer Jump. However, the larger, cooler, lower ionized zones that still contribute to the free bound continuum would not be captured by small slits and large discrepancies could be avoided. Our spectroscopic apertures, which have sizes of several kpc, would capture the full extent of the nebulae where temperatures are genuinely lower than in the small sub-regions that dominate the [O iii] emission.
6.3 Metal Abundance Patterns